# Many-Worlds Inverse Rendering

### v1

@@ Abstract.
Our work presents another solution: instead of differentiating a tentative surface locally, we differentiate a volumetric perturbation of a surface. We refer this as a “many-worlds“ representation because it models a non-interacting superposition of conflicting explanations (“worlds”) of the input dataset. Each world is optically isolated from others, leading to a new transport law that distinguishes our method from prior work based on exponential random media.
@@ 1. Introduction
Most recent work on this problem builds on the notion of emissive volumes [NeRF; GaussianSplats]. They are popular because the direct relationship between emission and observations leads to exceptionally well-behaved optimization problems.
@@
(2) $\vskip-1.42262pt\pi_{i+1}\leftarrow\pi_{i}-\eta\nabla_{\pi}\mathcal{L}(\mathcal{R}(\pi_{i})),\vskip-1.42262pt$
@@
Seen from a high level, this isn’t a new idea: for example, numerous prior works have observed how surface reconstruction problems become tractable once those surfaces have been “blurred” into a volume. A recent paper by Miller et al. [Miller2023Theory] interprets microflake volumes as the relaxation of a stochastic surface model and discusses a large body of related work that optimizes volumes to eventually obtain surfaces. While the extension from surfaces to exponential volumes does indeed address many problems, we argue that it is unnecessarily complex:
@@
    Extraction. Turning fuzzy inter-reflecting volumes into optically equivalent flattened surfaces presents its own challenges. Prior work related the bulk properties of opaque microfacets and microflakes [dupuy2016additional; Miller2023Theory], but this insight does not generalize to the richness of reflectance models in use in today’s rendering systems. In general, a complete pipeline will require a post-process optimization step to find a surface that locally matches a previously reconstructed volume. This breaks the elegance of simply discarding extra dimensions once the optimization converges (Equation 4).
@@ 2.1. Physically-based inverse rendering
Estimating the missing derivative requires sampling rays along shape boundaries (i.e., silhouettes). Prior work [li2018differentiable; Zhang2020PSDR; yan2022efficient; Zhang2023Projective] has extensively studied the theory and practice of estimating this boundary term. Another approach employs reparameterizations to convert the boundary contribution to neighboring regions, enabling simpler area-based formulations. These methods [Loubet2019Reparameterizing; bangaru2020unbiased] often require tracing auxiliary rays to detect the presence of neighboring silhouettes. For geometry representations based on signed distance fields (SDFs), specialized algorithms can alleviate some of this cost [vicini2022differentiable; Bangaru2022NeuralSDFReparam].
@@
The derivative of the rendering process can be decomposed into two types: the continuous part and the discontinuous part. We refer readers to Figure 4 in [Zhang2023Projective] for an illustration. Since our paper focuses on geometry optimization, we refer to the continuous part as the shading derivative and the discontinuous part as the boundary derivative in the following.
The shading derivative is defined everywhere on the surface and primarily affects the surface through its normal. Because shading depends on lighting, the smoothness of the illumination environment can have a marked effect on the robustness of optimization. Adjoint rendering methods [NimierDavid2020Radiative; Vicini2021PathReplay] exploit symmetries to compute the shading derivative with linear time complexity. Zeltner2021MonteCarlo analyze different strategies for variance reduction in this process. The boundary derivative, on the other hand, is only defined on the visibility silhouette of the surface. These surface derivatives are sparse in two senses:
@@
    A single view only observes a limited subset of all possible silhouettes, and a derivative step is likely to introduce a kink at this subset. Nicolet2021Large reparameterize meshes on a space that promotes smoothness to mitigate this issue.
@@
    The derivatives are only defined on the surface, not throughout the entire 3D space. Although light transport is simulated in this larger space, only the neighborhood of a surface can use the computed gradients. Regions far from the initial guess, must wait until the surface extends to them to receive gradients. mehta2023theory propose an elegant solution for vector graphics in 2D; however, they do not extend their approach to define surface derivatives in the entire 3D space.
@@
There is a substantial body of literature on approximating the PBR process to make the optimization more tractable. These methods often bypass the discontinuity problem by assuming knowledge of a shape mask (implicitly assuming the shape is directly observable), or by using approximated boundary derivative estimation algorithms like Nvdiffrast [Laine2020diffrast]. Additionally, most methods approximate the rendering equation for efficiency [physg2021; jin2023tensoir; zhang2021nerfactor; boss2021nerd]. These methods are physically-inspired and can handle complex scenes well, while our theory focuses on differentiating a fully physically-based rendering process.
NimierDavid2022Unbiased demonstrate that a physically-based non-emissive volume can be optimized to fit the input images. These methods are more physically accurate than the emissive volume methods but do not yield a surface representation in the end.
@@ 2.2. Emissive volume inverse rendering
A large body of work based on variants of the NeRF [NeRF] technique replaces the expensive rendering process with an emissive volume, which leads to convergent optimizations with excellent generalization to novel viewpoints. However, the final reconstruction is fog-like volume that does not accurately represent a solid surface.
Later work [Yariv2021Volume; wang2021neus; li2023neuralangelo; Miller2023Theory] added a surface prior to the volume model by introducing an SDF and deriving the volume density from it. Intuitively, volume particles should concentrate near the surface and decrease as one moves away from it. To model this effect, a symmetric bell-shaped distribution is added to the SDF value (centered at the zero-level set) to convert it into a volume density. The variance of the distribution is learnable during optimization: it begins with a high-variance distribution, mimicking the NeRF model, and gradually reduces to ensure that a surface can confidently be extracted in the end.
These SDF-based volume reconstructions can be seen as a blurry version of the surface-based methods without indirect effects. When the variance is zero, these methods simplify to surface-based methods [^1] This equivalence is theoretical. In practice, algorithms designed for volume optimization cannot handle such a Dirac-delta density distribution due to discontinuities.. With moderate variance, the effect of a solid surface, which we can interpret as “terminating a ray at a position with probability 11”, is now diffused to the neighborhood, interpretable as “terminating a ray in this neighborhood according to a probability density”. The blur eliminates the need to compute boundary derivatives.
@@ 2.3. Surfaces versus volumes
Several related works studied the relationship of surface and volume rendering: Heitz et al. [heitz2016multiple] and Dupuy et al. [dupuy2016additional] describe microfacet surfaces within the framework of random volumes. Vicini et al. [vicini2021non] incorporated non-exponential transmittance to account for correlations arising from opaque surfaces. Miller et al. [Miller2023Theory] formalize microflake volumes as the relaxation of a stochastic surface model and solve inverse problems using this representation. Seyb et al. [seyb2024microfacets] recently incorporated a richer set of spatial correlations, producing a forward rendering framework that spans the full continuum ranging from pure volume to surface-like interactions.
@@ 3.1. Motivation
Miller et al. [Miller2023Theory] recently studied the relationship between random surface distributions and microflake volumes. They showed that under reasonable assumptions, the distribution of surfaces can be aggregated into a 3D occupancy
@@
Our method departs from prior work in the following notable way: whereas existing inverse rendering methods have largely differentiated known surface-based or volume-based rendering algorithms, our work proposes a dedicated derivative transport law that operates within the derivative propagation step of radiative backpropagation [NimierDavid2020Radiative]. This means that the resulting derivatives do not correspond to the differentiation of any physically-based method. Indeed, a direct attempt to retrofit them to primal rendering would produce nonsensical results that violate energy conversation. Section 3.7 finally revisits this point and shows how the many-worlds representation can be adapted to the setting of primal rendering.
@@ 3.3. Transport in a tenuous volume
which follows from $e^{-x}\approx 1-x+O(x^{2})$. Plugging this into the RTE (9) and discarding a second-order term ($\mu_{\mathrm{s}}\mu_{\mathrm{t}}\sim\varepsilon^{2}$) yields the approximation $L_{\mathrm{i}}(0)\approx L_{\mathrm{i}}^{\hskip-0.85358pt\varepsilon}(0)$, where the latter is defined as
(13) $L_{\mathrm{i}}^{\hskip-0.85358pt\varepsilon}(0)\!\coloneqq\!\!\!\int_{0}^{s}\!\!\!\left[\mu_{\mathrm{s}}(t)\,L_{\mathrm{s}}(t)+\mu_{\mathrm{a}}(t)\,L_{\mathrm{e}}(t)-\mu_{\mathrm{t}}(t)\,L_{\mathrm{o}}(s)\right]\mathrm{d}t+L_{o}(s),\!$
@@ 3.4. Completing the volume model
To complete this model, we must instill meaning into the terms $\mu_{\mathrm{t}}$, $\mu_{\mathrm{s}}$, and $f_{p}$. Prior work often did so using volumetric analogs of microfacet surface models known as microflakes [Jakob2010Microflake]. Applications include optimization, volumetric level of detail, energy-conserving random walks, and neural fields [heitz2015sggx; heitz2016multiple; vicini2021non; Loubet2018Microflake; Zhang2023NeMF].
It is worth noting that the theory of microflakes actually builds on a more general notion of anisotropic radiative transport [shultis1988radiative; Jakob2010Microflake], which adopts directionally varying extinction and scattering coefficients arising from a random distribution of oriented particles:
@@
(17) $\mu_{\mathrm{t}}(\mathbf{x},\bm{\omega})$ $=\begin{cases}\varepsilon\,\alpha(\mathbf{x}),&\mathrm{if}\ \bm{\omega}\cdot\bm{\beta}(x)>0,\\ 0,&\mathrm{otherwise},\end{cases}$
@@
(21) $L_{\mathrm{i}}^{\hskip-0.85358pt\varepsilon}(0)=\varepsilon\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\mathrm{d}t+L_{o}(s),$
@@ 3.5. Many-worlds transport
The term $\mathcal{R}(\bar{S}+\varepsilon\cdot V_{S})$ in the numerator led to the derivation of the linearized transport law (21), which produces this perturbed image. The effect of subsequently subtracting $\mathcal{R}(\bar{S})$, dividing by $\varepsilon$, and then adding back $\mathcal{R}(\bar{S})$ is to cancel out the $\varepsilon$ scale factor so that the final transport law has a unit amount (thus $L_{\mathrm{i}}^{\hskip-0.85358pt1}$) of the many-worlds perturbation:
(22) $L_{\mathrm{i}}^{\hskip-0.85358pt1}(0)=\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\,\mathrm{d}t+L_{o}(s).$
@@ 3.6. Many-worlds derivative transport
$\partial_{\pi}L_{\mathrm{i}}^{\hskip-0.85358pt1}(0)$ $=\partial_{\pi}\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\,\mathrm{d}t+\underbracket{\partial_{\pi}L_{\mathrm{o}}(s)}_{=\ 0}$
@@
Figure 5. Many-worlds derivative transport. The propagated derivative at a sampled position $t$ is only weighted by the local color difference and the local occupancy respectively (Equation 23). In contrast to exponential or non-exponential (e.g., linear [vicini2021non]) volumes, the notion of transmittance disappears.
@@ 3.7. Primal rendering
The value $A$ specifies to what degree the volume masks the mean surface. This masking makes sense when $A\!\in\![0,1]$. But $A$ can be as large as $s$, which potentially exceeds $1$ by a significant amount, and this can cause $L_{\mathrm{i}}^{\hskip-0.85358pt1}(0)$ to take on unnaturally large or even negative values.
@@
${\hat{L}}_{\mathrm{i}}^{\hskip-0.85358pt1}(0)$ $=\frac{1}{s}\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\,\mathrm{d}t+L_{o}(s)$
@@
(26) ${\hat{L}}_{\mathrm{i}}^{\hskip-0.85358pt1}(0)=E\left[L_{\mathrm{o}}(X)\right]$
@@
The $1/s$ scaling is a natural design choice with a clear interpretation, but it is not unique. Other choices could also be used, as the rendering step is decoupled from the proposed derivative propagation [NimierDavid2020Radiative].
@@ 4. Discussion
Prior works on differentiable rendering of surfaces have proposed local formulations [Zhang2020PSDR [Zhang2020PSDR], Zhang2023Projective [Zhang2023Projective]] to express derivatives arising from the discontinuous visibility function. These local formulations provide a different viewpoint that can unlock significant efficiency gains. For example, consider a differentiable rendering of a triangle mesh occluding another object. Normally, this would require sampling light paths that graze the silhouette of the occluder, and finding such paths can be computationally expensive. However, it is is equivalent and generally easier to start from the middle by randomly picking a polygonal edge on the occluder along with a tangential direction, and then bidirectionally expanding it into a complete light path. This is referred to as a local formulation because it expresses the gradient computation as a low-dimensional integral over the space where derivatives actually arise (e.g., at silhouette points causing an occlusion).
@@
This derivation grounds the many-worlds method in an established theory and we made the following observations:
@@
An important detail that must be addressed is how the many-worlds transport equation (Equation 27) should be sampled in a Monte Carlo renderer. Since we don’t know the expected magnitude of $L_{\mathrm{o}}(t)$, the default strategy will simply sample all other terms proportional to their known contribution. In other words, by drawing positions $t$ with density $\alpha(t)/A$ and the surface at $s$ with discrete probability 1A/s1-A/s. While this is a sensible choice for primal rendering, it leads to a chicken-and-egg problem during optimization. It would imply that regions with $\alpha(\mathbf{x})\!\approx\!0$ are essentially never sampled, but we must clearly visit them sometimes to even consider the possibility of placing a surface there.
@@
(28) $\partial_{\pi}L_{\mathrm{i}}^{\hskip-0.85358pt1}(0)=\int_{0}^{s}\big(\underbracket{\partial_{\pi}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]}_{\text{(i)}}+\underbracket{\alpha^{+}(t)\partial_{\pi}L_{\mathrm{o}}(t)}_{\text{(ii)}}\big)\,\mathrm{d}t$
@@
This isn’t a new discovery: Nimier-David et al. [NimierDavid2022Unbiased] discuss the same issue in exponential volumes and address it by introducing generalizations of volume trackers that sample proportional to pure transmittance $T(t)$ instead of the default extinction-weighted transmittance $\mu_{\mathrm{t}}(t)T(t)$. Many-worlds transport has no transmittance, so the solution ends up being even simpler: a uniform sampling strategy suffices.

### v2

@@ Abstract.
Our work presents another solution: instead of differentiating a tentative surface locally, we differentiate a volumetric perturbation of a surface. We refer to this as a “many-worlds“ representation because it models a non-interacting superposition of conflicting explanations (“worlds”) of the input dataset. Each world is optically isolated from others, leading to a new transport law that distinguishes our method from prior work based on exponential random media.
@@ 1. Introduction
Most recent work on this problem builds on the notion of emissive volumes [Mildenhall et al., 2021; Kerbl et al., 2023]. They are popular because the direct relationship between emission and observations leads to exceptionally well-behaved optimization problems.
@@
(2) $\vspace{-.5mm}\pi_{i+1}\leftarrow\pi_{i}-\eta\nabla_{\pi}\mathcal{L}(\mathcal{R}(\pi_{i})),\vspace{-.5mm}$
@@
Seen from a high level, this isn’t a new idea: for example, numerous prior works have observed how surface reconstruction problems become tractable once those surfaces have been “blurred” into a volume. A recent paper by Miller et al. [2024] interprets microflake volumes as the relaxation of a stochastic surface model and discusses a large body of related work that optimizes volumes to eventually obtain surfaces. While the extension from surfaces to exponential volumes does indeed address many problems, we argue that it is unnecessarily complex:
@@
    Extraction. Turning fuzzy inter-reflecting volumes into optically equivalent flattened surfaces presents its own challenges. Prior work related the bulk properties of opaque microfacets and microflakes [Dupuy et al., 2016; Miller et al., 2024], but this insight does not generalize to the richness of reflectance models in use in today’s rendering systems. In general, a complete pipeline will require a post-process optimization step to find a surface that locally matches a previously reconstructed volume. This breaks the elegance of simply discarding extra dimensions once the optimization converges (Equation 4).
@@ 2.1. Physically-based inverse rendering
Estimating the missing derivative requires sampling rays along shape boundaries (i.e., silhouettes). Prior work [Li et al., 2018; Zhang et al., 2020; Yan et al., 2022; Zhang et al., 2023a] has extensively studied the theory and practice of estimating this boundary term. Another approach employs reparameterizations to convert the boundary contribution to neighboring regions, enabling simpler area-based formulations. These methods [Loubet et al., 2019; Bangaru et al., 2020] often require tracing auxiliary rays to detect the presence of neighboring silhouettes. For geometry representations based on signed distance fields (SDFs), specialized algorithms can alleviate some of this cost [Vicini et al., 2022; Bangaru et al., 2022].
@@
The derivative of the rendering process can be decomposed into two types: the continuous part and the discontinuous part. We refer readers to Figure 4 in [Zhang et al., 2023a] for an illustration. Since our paper focuses on geometry optimization, we refer to the continuous part as the shading derivative and the discontinuous part as the boundary derivative in the following.
The shading derivative is defined everywhere on the surface and primarily affects the surface through its normal. Because shading depends on lighting, the smoothness of the illumination environment can have a marked effect on the robustness of optimization. Adjoint rendering methods [Nimier-David et al., 2020; Vicini et al., 2021b] exploit symmetries to compute the shading derivative with linear time complexity. Zeltner et al. [2021] analyze different strategies for variance reduction in this process. The boundary derivative, on the other hand, is only defined on the visibility silhouette of the surface. These surface derivatives are sparse in two senses:
@@
    A single view only observes a limited subset of all possible silhouettes, and a derivative step is likely to introduce a kink at this subset. Nicolet et al. [2021] reparameterize meshes on a space that promotes smoothness to mitigate this issue.
@@
    The derivatives are only defined on the surface, not throughout the entire 3D space. Although light transport is simulated in this larger space, only the neighborhood of a surface can use the computed gradients. Regions far from the initial guess, must wait until the surface extends to them to receive gradients. Mehta et al. [2023] propose an elegant solution for vector graphics in 2D; however, they do not extend their approach to define surface derivatives in the entire 3D space.
@@
There is a substantial body of literature on approximating the PBR process to make the optimization more tractable. These methods often bypass the discontinuity problem by assuming knowledge of a shape mask (implicitly assuming the shape is directly observable), or by using approximated boundary derivative estimation algorithms like Nvdiffrast [Laine et al., 2020]. Additionally, most methods approximate the rendering equation for efficiency [Zhang et al., 2021a; Jin et al., 2023; Zhang et al., 2021b; Boss et al., 2021]. These methods are physically-inspired and can handle complex scenes well, while our theory focuses on differentiating a fully physically-based rendering process.
Nimier-David et al. [2022] demonstrate that a physically-based non-emissive volume can be optimized to fit the input images. These methods are more physically accurate than the emissive volume methods but do not yield a surface representation in the end.
@@ 2.2. Emissive volume inverse rendering
A large body of work based on variants of the NeRF [Mildenhall et al., 2021] technique replaces the expensive rendering process with an emissive volume, which leads to convergent optimizations with excellent generalization to novel viewpoints. However, the final reconstruction is fog-like volume that does not accurately represent a solid surface.
Later work [Yariv et al., 2021; Wang et al., 2021; Li et al., 2023; Miller et al., 2024] added a surface prior to the volume model by introducing an SDF and deriving the volume density from it. Intuitively, volume particles should concentrate near the surface and decrease as one moves away from it. To model this effect, a symmetric bell-shaped distribution is added to the SDF value (centered at the zero-level set) to convert it into a volume density. The variance of the distribution is learnable during optimization: it begins with a high-variance distribution, mimicking the NeRF model, and gradually reduces to ensure that a surface can confidently be extracted in the end.
These SDF-based volume reconstructions can be seen as a blurry version of the surface-based methods without indirect effects. When the variance is zero, these methods simplify to surface-based methods [^1] This equivalence is theoretical. In practice, algorithms designed for volume optimization cannot handle such a Dirac-delta density distribution due to discontinuities.. With moderate variance, the effect of a solid surface, which we can interpret as “terminating a ray at a position with probability $1$”, is now diffused to the neighborhood, interpretable as “terminating a ray in this neighborhood according to a probability density”. The blur eliminates the need to compute boundary derivatives.
@@ 2.3. Surfaces versus volumes
Several related works studied the relationship of surface and volume rendering: Heitz et al. [2016] and Dupuy et al. [2016] describe microfacet surfaces within the framework of random volumes. Vicini et al. [2021a] incorporated non-exponential transmittance to account for correlations arising from opaque surfaces. Miller et al. [2024] formalize microflake volumes as the relaxation of a stochastic surface model and solve inverse problems using this representation. Seyb et al. [2024] recently incorporated a richer set of spatial correlations, producing a forward rendering framework that spans the full continuum ranging from pure volume to surface-like interactions.
@@ 3.1. Motivation
Miller et al. [2024] recently studied the relationship between random surface distributions and microflake volumes. They showed that under reasonable assumptions, the distribution of surfaces can be aggregated into a 3D occupancy
@@
Our method departs from prior work in the following notable way: whereas existing inverse rendering methods have largely differentiated known surface-based or volume-based rendering algorithms, our work proposes a dedicated derivative transport law that operates within the derivative propagation step of radiative backpropagation [Nimier-David et al., 2020]. This means that the resulting derivatives do not correspond to the differentiation of any physically-based method. Indeed, a direct attempt to retrofit them to primal rendering would produce nonsensical results that violate energy conversation. Section 3.7 finally revisits this point and shows how the many-worlds representation can be adapted to the setting of primal rendering.
@@ 3.3. Transport in a tenuous volume
which follows from $e^{-x}\approx 1-x+O(x^{2})$. Plugging this into the RTE (9) and discarding a second-order term ($\mu_{\mathrm{s}}\mu_{\mathrm{t}}\sim\varepsilon^{2}$) yields the approximation $L_{\mathrm{i}}(0)\approx L_{\mathrm{i}}^{\varepsilon}(0)$, where the latter is defined as
(13) $L_{\mathrm{i}}^{\varepsilon}(0)\!\coloneqq\!\!\!\int_{0}^{s}\!\!\!\left[\mu_{\mathrm{s}}(t)\,L_{\mathrm{s}}(t)+\mu_{\mathrm{a}}(t)\,L_{\mathrm{e}}(t)-\mu_{\mathrm{t}}(t)\,L_{\mathrm{o}}(s)\right]\mathrm{d}t+L_{o}(s),\!$
@@ 3.4. Completing the volume model
To complete this model, we must instill meaning into the terms $\mu_{\mathrm{t}}$, $\mu_{\mathrm{s}}$, and $f_{p}$. Prior work often did so using volumetric analogs of microfacet surface models known as microflakes [Jakob et al., 2010]. Applications include optimization, volumetric level of detail, energy-conserving random walks, and neural fields [Heitz et al., 2015, 2016; Vicini et al., 2021a; Loubet and Neyret, 2018; Zhang et al., 2023b].
It is worth noting that the theory of microflakes actually builds on a more general notion of anisotropic radiative transport [Shultis and Myneni, 1988; Jakob et al., 2010], which adopts directionally varying extinction and scattering coefficients arising from a random distribution of oriented particles:
@@
(17) $\mu_{\mathrm{t}}(\mathbf{x},\bm{\omega})$ =$=\begin{cases}\varepsilon\,\alpha(\mathbf{x}),&\mathrm{if}\ \bm{\omega}\cdot\bm{\beta}(x)>0,\\ 0,&\mathrm{otherwise},\end{cases}$
@@
(21) $L_{\mathrm{i}}^{\varepsilon}(0)=\varepsilon\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\mathrm{d}t+L_{o}(s),$
@@ 3.5. Many-worlds transport
The term $\mathcal{R}(\bar{S}+\varepsilon\cdot V_{S})$ in the numerator led to the derivation of the linearized transport law (21), which produces this perturbed image. The effect of subsequently subtracting $\mathcal{R}(\bar{S})$, dividing by $\varepsilon$, and then adding back $\mathcal{R}(\bar{S})$ is to cancel out the $\varepsilon$ scale factor so that the final transport law has a unit amount (thus $L_{\mathrm{i}}^{1}$) of the many-worlds perturbation:
(22) $L_{\mathrm{i}}^{1}(0)=\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\,\mathrm{d}t+L_{o}(s).$
@@ 3.6. Many-worlds derivative transport
[…]
@@
Figure 5. Many-worlds derivative transport. The propagated derivative at a sampled position $t$ is only weighted by the local color difference and the local occupancy respectively (Equation 23). In contrast to exponential or non-exponential (e.g., linear [Vicini et al., 2021a]) volumes, the notion of transmittance disappears.
@@ 3.7. Primal rendering
The value $A$ specifies to what degree the volume masks the mean surface. This masking makes sense when $A\!\in\![0,1]$. But $A$ can be as large as $s$, which potentially exceeds $1$ by a significant amount, and this can cause $L_{\mathrm{i}}^{1}(0)$ to take on unnaturally large or even negative values.
@@
${\hat{L}}_{\mathrm{i}}^{1}(0)$ $=\frac{1}{s}\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\,\mathrm{d}t+L_{o}(s)$
@@
(26) ${\hat{L}}_{\mathrm{i}}^{1}(0)=E\left[L_{\mathrm{o}}(X)\right]$
@@
The $1/s$ scaling is a natural design choice with a clear interpretation, but it is not unique. Other choices could also be used, as the rendering step is decoupled from the proposed derivative propagation [Nimier-David et al., 2020].
@@ 4. Discussion
Prior works on differentiable rendering of surfaces have proposed local formulations [Zhang et al. [2020], Zhang et al. [2023a]] to express derivatives arising from the discontinuous visibility function. These local formulations provide a different viewpoint that can unlock significant efficiency gains. For example, consider a differentiable rendering of a triangle mesh occluding another object. Normally, this would require sampling light paths that graze the silhouette of the occluder, and finding such paths can be computationally expensive. However, it is is equivalent and generally easier to start from the middle by randomly picking a polygonal edge on the occluder along with a tangential direction, and then bidirectionally expanding it into a complete light path. This is referred to as a local formulation because it expresses the gradient computation as a low-dimensional integral over the space where derivatives actually arise (e.g., at silhouette points causing an occlusion).
@@
This derivation grounds the many-worlds method in an established theory and we make the following observations:
@@
An important detail that must be addressed is how the many-worlds transport equation (Equation 27) should be sampled in a Monte Carlo renderer. Since we don’t know the expected magnitude of $L_{\mathrm{o}}(t)$, the default strategy will simply sample all other terms proportional to their known contribution. In other words, by drawing positions $t$ with density $\alpha(t)/A$ and the surface at $s$ with discrete probability $1-A/s$. While this is a sensible choice for primal rendering, it leads to a chicken-and-egg problem during optimization. It would imply that regions with $\alpha(\mathbf{x})\!\approx\!0$ are essentially never sampled, but we must clearly visit them sometimes to even consider the possibility of placing a surface there.
@@
(28) $\partial_{\pi}L_{\mathrm{i}}^{1}(0)=\int_{0}^{s}\big{(}\underbracket{\partial_{\pi}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]}_{\text{(i)}}+\underbracket{\alpha^{+}(t)\partial_{\pi}L_{\mathrm{o}}(t)}_{\text{(ii)}}\big{)}\,\mathrm{d}t$
@@
This isn’t a new discovery: Nimier-David et al. [2022] discuss the same issue in exponential volumes and address it by introducing generalizations of volume trackers that sample proportional to pure transmittance $T(t)$ instead of the default extinction-weighted transmittance $\mu_{\mathrm{t}}(t)T(t)$. Many-worlds transport has no transmittance, so the solution ends up being even simpler: a uniform sampling strategy suffices.